OpenStudy (anonymous):

Cos(x+y)+cos(x-y)=2cosxcosy

OpenStudy (anonymous):

\[\cos(x+y)=\cos x\cos y-\sin x\sin y\\ \cos(x-y)=\cos x\cos y+\sin x\sin y\]What happens when you add?